# General relativity and the double-zero eigenvalue

### v1

@@ 6 Discussion
A last problem we would like to briefly discuss is related to the question of the structure and nature of the theory that contains all versal unfoldings starting from a given one. We have begun to answering this question for general relativity, but one could perhaps apply similar ideas to other areas of fundamental theoretical physics. What is the nature of the versal families that describe the unfoldings of the Einstein, Maxwell, Schrödinger, or Dirac equations? A way to begin to answer this question may be by looking at systems that couple gravity to other fields in a specialized context, for instance, in a isotropic universe, or a spherically symmetric star, etc. In this sense, one arrives at a new ‘theory’ that ‘pierces’ through the standard ones at degenerate points (e.g., general relativity at the ‘point’ (5.5)) and contains all such versal final forms of the situation one started with. For example, the landscape construction of string theory may have something to do with the versal unfoldings that the original theory contains and these in turn may lead to a new classification of possible string solutions. Another example of this sort would be to start with say the Einstein-Maxwell equations in some simplified context (say for a simple Bianchi model) and unfold. What would be the nature of the resulting theory? We believe that there is a lot to learn by deploying this program for the versal unfoldings of a given theory and discovering their relations to those of a suitably chosen ‘neighbor’.
@@ Acknowledgments
The author is grateful to Ignatios Antoniadis and Alexander Yefremov for useful discussions and various aspects of this work. This research was funded by RUDN University, scientific project number FSSF-2023-0003.

### v2

@@ 4 Discussion
A last problem we would like to briefly discuss is related to the question of the structure and nature of the theory that contains all versal unfoldings starting from a given one. We have begun to answering this question for general relativity, but one could perhaps apply similar ideas to other areas of fundamental theoretical physics. What is the nature of the versal families that describe the unfoldings of the Einstein, Maxwell, Schrödinger, or Dirac equations? A way to begin to answer this question may be by looking at systems that couple gravity to such fields in a specialized context, for instance, in a isotropic universe, or a spherically symmetric star, etc. In this sense, one arrives at a new ‘theory’ that ‘pierces’ through the standard ones at degenerate points (e.g., general relativity at the ‘point’ (3.15)), and contains all such versal final forms of the situation one started with. For example, the landscape construction of string theory may have something to do with the versal unfoldings that the original theory contains and these in turn may lead to a new classification of possible string solutions. Another example of this sort would be to start with say the Einstein-Maxwell equations in some simplified context (say for a simple Bianchi model) and unfold. What would be the nature of the resulting theory? We believe that there is a lot to learn by deploying this program for the versal unfoldings of a given theory and discovering their relations to those of a suitably chosen ‘neighbor’.
@@ Acknowledgments
The author is grateful to three anonymous referees whose suggestions and constructive comments have led to a better version of this work, and to Ignatios Antoniadis and Alexander Yefremov for useful discussions related to the work presented in this paper. This research was funded by RUDN University, scientific project number FSSF-2023-0003.